# Interpretable and Equation-Free Response Theory for Complex Systems

### v1

@@ Abstract
Response theory provides a pathway for understanding the sensitivity of a system and, more in general, to predict how its statistical properties change as a possibly time-dependent perturbation is applied. Recently discovered general forms of the celebrated Fluctuation-Dissipation Theorem allow for expressing response operators as correlation functions of suitably defined observables in the unperturbed state, also when such a state is far from equilibrium. In the case of complex and multiscale systems, to achieved enhanced practical applicability, response theory must be interpretable, capable of focusing of relevant timescales, and amenable to implemented by data-driven approaches that are potentially equation-agnostic. Complex systems typically exhibit a hierarchy of temporal behaviors, and unresolved or undesired timescales can obscure the dominant mechanisms driving macroscopic responses. As an element of this desired framework, in the spirit of Markov state modelling, we propose here a comprehensive analysis of the linear and nonlinear response of Markov chains to general time-dependent perturbations. We obtain simple and easily implementable formulas that can be used to predict the response of observables as well as higher-order correlations of the system. The methodology proposed here can be implemented in a purely data-driven setting and even if we do not know the underlying evolution equations. The use of algebraic expansions inspired by Koopmanism allow to elucidate the role of different time scales and find explicit and interpretable expressions for the Green’s functions at all orders. We illustrate our methodology in a very simple yet instructive metastable system. Finally, our results provide a dynamical foundation for the Prony method, which is commonly used for the statistical analysis of discrete time signals.
@@ I Introduction
Response theory in statistical mechanics constitutes a powerful framework for analyzing the behavior of a large variety of systems subjected to external perturbations. It provides a powerful and unifying paradigm for connecting the microscopic dynamics and reference statistical properties of a system to its macroscopic response under external influences. Its foundations have been widely discussed in the mathematical literature and its applications permeate various domains of physics, chemistry, biology, materials science, and quantitative social sciences Hänggi and Thomas (1982); Baiesi and Maes (2013); Sarracino and Vulpiani (2019). At the core of response theory lies the formulation of response functions, which quantify the system’s reaction to acting forcings. In the case of systems at thermodynamic equilibrium and considering for the moment only the linear approximation to the response, the fluctuation-dissipation theorem (FDT) establishes that such functions are expressed in terms of time-lagged correlation between suitably defined observables on the unperturbed state Kubo (1966).
@@
However, response theory is not limited to equilibrium systems Hairer and Majda (2010). For systems driven far from equilibrium possessing smooth invariant measure with respect to Lebesgue - which is realized in the case of stochastic dynamical systems forced by sufficiently non-degenerate noise - it is possible to express the response formulas in terms of more cumbersome yet well-defined time-lagged correlations, which implies the existence of a clear correspondence between forced and free fluctuations of the system Marconi et al. (2008); Pavliotis (2014). Things become more problematic in the case of nonequilibrium systems described by deterministic dynamics featuring contraction of the phase space. Here, as a result of the singularity of the invariant measure with respect to Lebesgue, the FDT does not apply and there is no full equivalence between forced and free fluctuations. Yet, making suitable assumptions on the dynamics it is possible to establish a response theory also in this case. The original results proposed by Ruelle, which required fairly restructive conditions of uniform hyperbolicity Ruelle (1998a, 2009), have then been clarified and extended using functional analytical techniques Liverani and Gouëzel (2006); Baladi (2014); Baladi et al. (2017).
@@
When the perturbation is large or when the system exhibits a very amplified response, the linear approximation linking the amplitudes of the forcing and of the response breaks down. Extending response theory to the nonlinear case involves considering higher-order terms in the system’s response and exploring how these terms contribute to the system’s behavior Ruelle (1998b); Lucarini (2008, 2009). Nonlinear response operators have a rather convoluted structure and depend on multiple time variables. Looking at the nonlinear response amounts to exploring more complex interplay between internal feedbacks of the system and acting forcings and, if more than one forcing is present, allows for understanding the interplay - which can be synergistic or antagonistic - of the various acting forcings. Nonlinear effects become quantitatively dominant in the proximity of critical transitions, which are associated with the divergence of the response of the system. However, the occurrence of such divergent behavior can be captured simply by looking at the linear response of the system Chekroun et al. (2020); Santos Gutiérrez and Lucarini (2022); Lucarini and Chekroun (2023a).
@@
A key difficulty of response formulas is that they are based on expressions that do not provide a clear imprint of the dominant modes of variability of the system. A way forward in this direction is provided by Koopmanism Budišić et al. (2012); Kutz et al. (2016), which, roughly speaking, transforms nonlinear dynamics into a linear framework in an infinite dimensional space of observables, where the information is contained in the Koopman eigenfunctions and modes. Metastable regions of the system, where the system can identified by studying the level sets of the dominant modes of the Koopman operator Froyland et al. (2014). In practice, the Koopman operator is often approximated using data-driven techniques such as the extended DMD Brunton and Kutz (2022).
@@
By linking response theory with Koopmanism, one can derive interpretable representations of the system’s response to perturbations, The use of Koopmanism radically improves the predictive power of the FDT because it enables the decomposition of the response into contributions from distinct modes of natural variability of the system Santos Gutiérrez and Lucarini (2022); Lucarini and Chekroun (2023a). Recently, we have being successful in merging algorithmically response theory and Koopmanism Zagli et al. (2024) and in showing that Koopmanism provides a pathway for extending response theory to the very relevant yet so far unexplored case where the stochasticity includes jump processes Chekroun et al. (2024).x While these preliminary results are extremely encouraging both in theoretical and practical terms, a nontrivial hurdle that still needs to be overcome is the applicability of this methodology to very high-dimensional systems.
@@ I.1 A Possible Pipeline for an Interpretable and Equation-free Response Theory
A Markov chain is a stochastic process that describes transitions between a finite set of states according to a probability distribution given by the so-called Markov matrix, where the future state depends only on the current state and not on past states. Markov chains can be associated with directed graphs, whereby the states correspond to the nodes of the network, and the entries of the Markov matrix give the weight of the directed edge between two nodes Norris (1998); Behrends (2000). Markov chains are extremely relevant to understand the properties of general dynamical systems Bowen (1970); Attal (2010). The so-called Ulam method Ulam (1960) approximates the Perron-Frobenius operator, which pushes forward the probability distributions of a dynamical system Baladi (2000), by a Markov matrix whose entries represent transition probabilities between partition elements occurring for finite time horizons. The Ulam Conjecture states that such finite-rank approximation the dominant eigenvalues and eigenfunctions of the finite-size approximations converge to those of the true operator in the limit of finer and finer partition, thus preserving eventual metastable structures Froyland (1998).
While the Ulam method is by itself essentially a brute-force approximation, and its convergence is usually slow Ding et al. (2002), it is possible to use it very effectively. Specifically, Markov state modeling (MSM) is a smart Ulam method that is particularly effective for studying systems with complex dynamics that evolve across multiple timescales. In this framework, the continuous state space of a system is (optimally - which is key here) discretized into a finite number of states, with transitions between these states governed by a Markov chain Bussi et al. (2006); Pande et al. (2010); Bowman et al. (2014); Husic and Pande (2018). The derived Markov chain contains all the information needed to describe the statistics and dynamics of the system at the coarse-grained level. Metastable regions can be identified on the basis of the states identified by the algorithm. In this case, the slowest time scales are associated with the relaxation between the main metastable states, with modest or non-existent gap for the dynamics occurring within each metastable region. A maximally reduced version of MSM targets directly the metastable states and studies exclusively the transition rates between such states Bittracher et al. (2018).
@@
Comprehensive and easily implementable response formulas for Markov chains are - apart from their intrinsic interest - of great practical utility in the analysis of a complex system because they can be directly applied to its coarse-grained representation constructed according to MSM protocols and thus bypassing (and obliviously to) the underlying evolution equations. We have discussed elsewhere how the invariant measure of a Markov chain responds to a time-independent perturbation, presenting explicit bounds for the validity of classic perturbative approach and providing explicit formulas for linear and nonlinear terms, including renormalized results obtained by infinite summation of all the perturbative terms Lucarini (2016); Santos Gutiérrez and Lucarini (2020). Independent results that delve more deeply in the physical interpretation of the response formulas have been recently reported Aslyamov and Esposito (2024a, b).
@@
Treating carefully the transition from the microscopic description of a system to its heavily coarse-grained representation as a discrete Markov process and linking the properties of fluctuations and response across the scale is an extremely challenging task, see the excellent review papers Zhang et al. (2012); Falasco and Esposito (2025). Here we set ourselves in a much simplified setting. The scenario behind this paper is that we assume that upstream of our work someone has carefully constructed a coarse-grained representation of the system as a discrete Markov chain, e.g. by applying Markov state modeling to a complex, possibly multiscale system and has observed the system in a reference state and in a slightly perturbed state. Both states are characterised by autonomous dynamics. The goal we have is to predict how different time modulations of the forcing impacts the statistical properties of the coarse-grained system. Hence, by construction, we will neglect the subscale processes.
@@
We will investigate the interplay of coarse-graining and response and derive formulas that are able to predict the linear as well as higher order response of the coarse-grained system for general, time-dependent perturbation via simple matrix relations. Our ability to treat explicitly time-independent perturbation is, as far as we know, novel, and goes in the direction of analysis of entropy production for non autonomous systems Seifert (2005). Interestingly, response formulas can be derived for observables as well as for lagged correlations between observables, thus allowing for predicting how the forcing impacts the variability of the system. The latter had been attempted in a previous work but only in the case of static forcings Lucarini and Wouters (2017). We will also provide a simple but possibly very instructive novel way of expressing the linear and nonlinear response operators for Markov chains by taking advantage of the Koopman formalism for finite-state processes that clarifies the roles of the time scales that are intrinsic to the system. The derivation and discussion of response formulas is presented in Sect. II. In order to illustrate some of our findings, we will provide in Sect. IV a proof-of-concept application of some of our key results on a simple yet instructive two-dimensional (2D) Langevin equation closely related to an example provided in Klus et al. (2016) which is characterized by nontrivial metastability properties. In Sect. V we present a discussion of our results and perspectives for future investigations. Additionally App. A, App. B, and App. C provide general formulas for the nonlinear response of observables, the linea response of correlation functions, and the linear response of entropy production to time-dependent perturbations, respectively.
@@ II Response Theory for Markov Chains: Time-Dependent Perturbations
We set ourselves in the same framework described in Lucarini (2016); Santos Gutiérrez and Lucarini (2020). Let us consider a mixing $N$-state ($N$ is finite) Markov process defined by the matrix $\mathcal{M}\in\mathbb{R}^{N\times N}$. $\mathcal{M}_{ij}\geq 0$ is a stochastic matrix that measures the probability of reaching the state $i$ at time $n$ given that at time $n-1$ the system is in the state $j$. Since the process is mixing $\exists p\geq 1|\mathcal{M}^{p}_{ij}>0$. We consider the eigenvalue problem $\mathcal{M}\mathbf{v}=\lambda{v}$. For the Perron-Frobenius theorem, we have that $\exists!{\nu_{inv}}|\mathcal{M}{\nu_{inv}}={\nu_{inv}}$, so that ${\nu_{inv}}\in\mathbb{R}^{N\times 1}$ defines the invariant measure associated with unitary eigenvalue Seneta (1973). We also have pairs $\{\lambda_{j},{\nu_{j}}\}$, such that $\mathcal{M}{\nu_{j}}=\lambda_{j}{\nu_{j}}$, with $|\lambda_{j}|<1$ and $\nu_{j}\in\mathbb{R}^{N\times 1}$ for $j=2,\ldots N$. Additionally, we have that $\sum_{i=1}^{N}(\nu_{inv})_{i}=1$ and $\sum_{i=1}^{N}(\nu_{j})_{i}=0$, $j>1$.
Let’s now consider a perturbation of the form $\mathcal{M}\rightarrow\mathcal{M}_{\epsilon,n}=\mathcal{M}+\epsilon f(n)m$, where $f:\mathbb{N}\rightarrow\mathbb{R}$ defines a time-dependent modulation with $|f(n)|<1$, $m\in\mathbb{R}^{N\times N}$ , and $\epsilon$ is a small real number. We impose that $\mathcal{M}_{\epsilon,n}$ is at all times a stochastic matrix. Hence, $\sum_{i}m_{ij}=0$. The perturbed Markov chain evolves according to the following law:
@@
We plug $\nu(n)=\nu_{inv}+\epsilon\nu^{(1)}(n)+h.o.t.$ in the equation above and collect the terms proportional to $\epsilon$. The conditions behind the applicability of the perturbative approach are discussed in detail in Lucarini (2016); Santos Gutiérrez and Lucarini (2020) and will not be repeated here; see also Abbas et al. (2016) and recent review devoted to continuous time Markov chains Mitrophanov (2024). It suffices here to say that one requires the existence of a finite spectral gap for the operator $\mathcal{M}$. We then assume that $\epsilon$ is sufficiently small so that the perturbative expansions converge. We obtain:

### v2

@@ Abstract
Response theory provides a pathway for understanding the sensitivity of a system and, more generally, to predict how its statistical properties change as a possibly time-dependent perturbation is applied. Recently discovered general forms of the celebrated Fluctuation-Dissipation Theorem allow for expressing response operators as correlation functions of suitably defined observables in the unperturbed state, also when such a state is far from equilibrium. In the case of complex and multiscale systems, to achieved enhanced practical applicability, response theory must be interpretable, capable of focusing of relevant timescales, and amenable to implemented by data-driven approaches that are potentially equation-agnostic. Complex systems typically exhibit a hierarchy of temporal behaviors, and unresolved or undesired timescales can obscure the dominant mechanisms driving macroscopic responses. As an element of this desired framework, in the spirit of Markov state modelling, we propose here a comprehensive analysis of the linear and nonlinear response of Markov chains to general time-dependent perturbations. We obtain simple and easily implementable formulas that can be used to predict the response of observables as well as higher-order correlations of the system. The methodology proposed here can be implemented in a purely data-driven setting and even if we do not know the underlying evolution equations. The use of algebraic expansions inspired by Koopmanism allows to elucidate the role of different time scales and to find explicit and interpretable expressions for the Green’s functions at all orders. This is a major advantage of the framework proposed here. We illustrate our methodology in a very simple yet instructive metastable system. Finally, our results provide a dynamical foundation for the Prony method, which is commonly used for the statistical analysis of discrete time signals.
@@ I Introduction
Response theory in statistical mechanics constitutes a powerful framework for analyzing the behavior of a large variety of systems subjected to external perturbations. It provides a powerful and unifying paradigm for connecting the microscopic dynamics and reference statistical properties of a system to its macroscopic response under external influences. Its foundations have been widely discussed in the mathematical literature and its applications permeate various domains of physics, chemistry, biology, materials science, and quantitative social sciences [1, 2, 3]. At the core of response theory lies the formulation of response functions, which quantify the system’s reaction to acting forcings. In the case of systems at thermodynamic equilibrium and considering for the moment only the linear approximation to the response, the fluctuation-dissipation theorem (FDT) establishes that such functions are expressed in terms of time-lagged correlation between suitably defined observables on the unperturbed state [4].
@@
However, response theory is not limited to equilibrium systems [5]. For systems driven far from equilibrium possessing smooth invariant measure with respect to Lebesgue - which is realized in the case of stochastic dynamical systems forced by sufficiently non-degenerate noise - it is possible to express the response formulas in terms of more cumbersome yet well-defined time-lagged correlations, which implies the existence of a clear correspondence between forced and free fluctuations of the system [6, 7]. Things become more problematic in the case of nonequilibrium systems described by deterministic dynamics featuring contraction of the phase space. Here, as a result of the singularity of the invariant measure with respect to Lebesgue, the FDT does not apply and there is no full equivalence between forced and free fluctuations. Yet, making suitable assumptions on the dynamics it is possible to establish a response theory also in this case. The original results proposed by Ruelle, which required fairly restructive conditions of uniform hyperbolicity [8, 9], have then been clarified and extended using functional analytical techniques [10, 11, 12].
@@
When the perturbation is large or when the system exhibits a very amplified response, the linear approximation linking the amplitudes of the forcing and of the response breaks down. Extending response theory to the nonlinear case involves considering higher-order terms in the system’s response and exploring how these terms contribute to the system’s behavior [13, 14, 15]. Nonlinear response operators have a rather convoluted structure and depend on multiple time variables. Looking at the nonlinear response amounts to exploring more complex interplay between internal feedbacks of the system and acting forcings and, if more than one forcing is present, allows for understanding the interplay - which can be synergistic or antagonistic - of the various acting forcings. Nonlinear effects become quantitatively dominant in the proximity of critical transitions, which are associated with the divergence of the response of the system. However, the occurrence of such divergent behavior can be captured simply by looking at the linear response of the system [16, 17, 18].
@@
A key difficulty of response formulas is that they are based on expressions that do not provide a clear imprint of the dominant modes of variability of the system. A way forward in this direction is provided by Koopmanism [19, 20], which, roughly speaking, transforms nonlinear dynamics into a linear framework in an infinite dimensional space of observables, where the information is contained in the eigenfunction and modes of the Koopman (or Kolmogorov in the case of stochastic dynamics) operator. From now on, with an abuse of language, we will refer to the Koopman operator also in the stochastic case. Metastable regions of the system, where the system can identified by studying the level sets of the dominant modes of the Koopman operator [21]. In practice, the Koopman operator is often approximated using data-driven techniques such as the extended DMD [22]; see [23] for a comprehensive review of DMD methods.
@@
By linking response theory with Koopmanism, one can derive interpretable representations of the system’s response to perturbations, The use of Koopmanism radically improves the predictive power of the FDT because it enables, by linearity, the decomposition of the response into contributions from distinct modes of natural variability of the system [17, 18]. Recently, we have been successful in merging algorithmically response theory and Koopmanism [24] and in showing that Koopmanism provides a pathway for extending response theory to the very relevant yet so far unexplored case where the stochasticity includes jump processes [25]. While these preliminary results are extremely encouraging both in theoretical and practical terms, a nontrivial hurdle that still needs to be overcome is the applicability of this methodology to very high-dimensional systems. A possible way ahead might rely on the use of the recently proposed multiplicative DMD algorithm [26], whereby the dictionary used for approximating the Koopman operator is given by the characteristic functions of the cells of a suitably constructed Voronoï tessellation [27]; see further comments in Sect. V.
@@ I.1 A Possible Pipeline for an Interpretable and Equation-free Response Theory
A Markov chain is a stochastic process that describes transitions between a finite set of states according to a probability distribution given by the so-called Markov matrix, where the future state depends only on the current state and not on past states. Markov chains can be associated with directed graphs, whereby the states correspond to the nodes of the network, and the entries of the Markov matrix give the weight of the directed edge between two nodes [28, 29]. Markov chains are extremely relevant to understand the properties of general dynamical systems [30, 31]. The so-called Ulam method [32] approximates the Perron-Frobenius operator, which pushes forward the probability distributions of a dynamical system [33], by a Markov matrix whose entries represent transition probabilities between partition elements occurring for finite time horizons. The Ulam Conjecture states that such finite-rank approximation the dominant eigenvalues and eigenfunctions of the finite-size approximations converge to those of the true operator in the limit of finer and finer partition, thus preserving eventual metastable structures [34].
While the Ulam method is by itself essentially a brute-force approximation, and its convergence is usually slow [35], it is possible to use it very effectively. Specifically, Markov state modeling (MSM) is a smart Ulam method that is particularly effective for studying systems with complex dynamics that evolve across multiple timescales. In this framework, the continuous state space of a system is (optimally - which is key here) discretized into a finite number of states, with transitions between these states governed by a Markov chain [36, 37, 38, 39]. The derived Markov chain contains all the information needed to describe the statistics and dynamics of the system at the coarse-grained level. Metastable regions can be identified on the basis of the states identified by the algorithm. In this case, the slowest time scales are associated with the relaxation between the main metastable states, with modest or non-existent gap for the dynamics occurring within each metastable region. A maximally reduced version of MSM targets directly the metastable states and studies exclusively the transition rates between such states [40].
@@
Comprehensive and easily implementable response formulas for Markov chains are - apart from their intrinsic interest - of great practical utility in the analysis of a complex system because they can be directly applied to its coarse-grained representation constructed according to MSM protocols and thus bypassing (and obliviously to) the underlying evolution equations.
We discussed elsewhere how the invariant measure of a Markov chain responds to a time-independent perturbation, presenting explicit bounds for the validity of classic perturbative approach and providing explicit formulas for linear and nonlinear terms, including renormalized results obtained by infinite summation of all the perturbative terms [41, 42]. Independent results that delve more deeply in the physical interpretation of the response formulas have been recently reported [43, 44].
@@
Treating carefully the transition from the microscopic description of a system to its heavily coarse-grained representation as a discrete Markov process and linking the properties of fluctuations and response across the scale is an extremely challenging task, see the excellent review papers [45, 46]. Here we set ourselves in a much simplified setting. The scenario behind this paper is that we assume that upstream of our work someone has carefully constructed a coarse-grained representation of the system as a discrete Markov chain, e.g. by applying Markov state modeling to a complex, possibly multiscale system and has observed the system in a reference state and in a slightly perturbed state. Both states are characterised by autonomous dynamics. The goal we have is to predict how different time modulations of the forcing impacts the statistical properties of the coarse-grained system. Hence, by construction, we will neglect the subscale processes.
@@
We will investigate the interplay of coarse-graining and response and derive formulas that are able to predict the linear as well as higher order response of the coarse-grained system for general, time-dependent perturbation via simple matrix relations. Our ability to treat explicitly time-independent perturbation is, as far as we know, novel, and goes in the direction of analysis of entropy production for non autonomous systems [47]. Interestingly, response formulas can be derived for observables as well as for lagged correlations between observables, thus allowing for predicting how the forcing impacts the variability of the system. The latter had been attempted in a previous work but only in the case of static forcings [48]. We will also provide a simple but possibly very instructive novel way of expressing the linear and nonlinear response operators for Markov chains by taking advantage of the Koopman formalism for finite-state processes that clarifies the roles of the time scales that are intrinsic to the system. The derivation and discussion of response formulas is presented in Sect. II for observables in in Sect. III for correlations. In order to illustrate some of our findings, we will provide in Sect. IV a proof-of-concept application of some of our key results on a simple yet instructive two-dimensional (2D) Langevin equation closely related to an example provided in [49] which is characterized by nontrivial metastability properties. In Sect. V we present a discussion of our results and perspectives for future investigations. Additionally App. A, App. B, and App. C provide general formulas for the nonlinear response of observables, the linea response of correlation functions, and the linear response of entropy production to time-dependent perturbations, respectively.
@@ II Response Theory for Markov Chains: Time-Dependent Perturbations
We set ourselves in the same framework described in [41, 42]. Let us consider a mixing $N$-state ($N$ is finite) Markov process defined by the matrix $\mathcal{M}\in\mathbb{R}^{N\times N}$. $\mathcal{M}_{ij}\geq 0$ is a stochastic matrix that measures the probability of reaching the state $i$ at time $n$ given that at time $n-1$ the system is in the state $j$. Since the process is mixing we can reach any state $i$ starting from any state $j$ is we wait a sufficiently long time, or, more specifically $\exists p\geq 1|\mathcal{M}^{p}_{ij}>0$. We consider the eigenvalue problem $\mathcal{M}\mathbf{v}=\lambda{v}$. For the Perron-Frobenius theorem, there is a unique invariant measure, i.e. $\exists!{\nu_{inv}}|\mathcal{M}{\nu_{inv}}={\nu_{inv}}$, so that ${\nu_{inv}}\in\mathbb{R}^{N\times 1}$ defines the invariant measure associated with unitary eigenvalue [50]. We also have pairs $\{\lambda_{j},{\nu_{j}}\}$, such that $\mathcal{M}{\nu_{j}}=\lambda_{j}{\nu_{j}}$, with $|\lambda_{j}|<1$ and $\nu_{j}\in\mathbb{R}^{N\times 1}$ for $j=2,\ldots N$. Additionally, we have that $\sum_{i=1}^{N}(\nu_{inv})_{i}=1$ and $\sum_{i=1}^{N}(\nu_{j})_{i}=0$, $j>1$.
Let us now consider a perturbation of the form $\mathcal{M}\rightarrow\mathcal{M}_{\epsilon,n}=\mathcal{M}+\epsilon f(n)m$, where $f:\mathbb{N}\rightarrow\mathbb{R}$ defines a time-dependent modulation with $|f(n)|<1$, $m\in\mathbb{R}^{N\times N}$ , and $\epsilon$ is a small real number. We impose that $\mathcal{M}_{\epsilon,n}$ is at all times a stochastic matrix. Hence, $\sum_{i}m_{ij}=0$. The perturbed Markov chain evolves according to the following law:
@@
We plug $\nu(n)=\nu_{inv}+\epsilon\nu^{(1)}(n)+h.o.t.$ in the equation above and collect the terms proportional to $\epsilon$. The conditions behind the applicability of the perturbative approach are discussed in detail in [41, 42] and will not be repeated here; see also [51] and recent review devoted to continuous time Markov chains [52]. It suffices here to say that one requires the existence of a finite spectral gap for the operator $\mathcal{M}$. We then assume that $\epsilon$ is sufficiently small so that the perturbative expansions converge. We obtain: